In a wave graph. initial conditions were changed, modifying the relationship between HAMP and UAMP by applying different multiplication factors. investigate the effect and present the analysis.  The results are as follows:

UAMP= sqrt(g*h0)HAMP/h0

elevation 12.0 100.5 x (meters) elevation (m) 9.5 9.7 9.9 10.1 10.3 10.5

by applying multiplication factor : (0.5)

UAMP= sqrt(g*h0)HAMP/h0*0.5

elevatior 12.0 100.5 x (meters) elevation (m) 9.5 9.7 9.9 10.1 10.3 10.5

by applying multiplication factor : (0.75)

UAMP= sqrt(g*h0)HAMP/h0*0.75

elevation 12.0 100.5 x (meters) elevation (m) 9.5 9.7 9.9 10.1 10.3 10.5

by applying multiplication factor : (-0.5)

UAMP= sqrt(g*h0)HAMP/h0*(-0.5)

elevatior 12.0 100.5 x (meters) elevation (m) 9.5 9.7 9.9 10.1 10.3 10.5
Answer #1

(i) Using a formula, we have

UAMP = \sqrt{}(g h0) HAMP / h0

where, g = acceleration due to gravity = 9.8 m/s2

h0 = initial height = 9.5 m

HAMP = final height = 10.5 m

then, we get

UAMP = \sqrt{}[(9.8 m/s2) (9.5 m) [10.5 m)] / (9.5 m)

UAMP = \sqrt{}102.9 m2/s2

UAMP = 10.1 m/s

(ii) By applying a multiplication factor, then we get

UAMP = [(10.1 m/s) (0.5)]

UAMP = 5.05 m/s

(iii) By applying a multiplication factor, then we get

UAMP = [(10.1 m/s) (0.75)]

UAMP = 7.57 m/s

(iv) By applying a multiplication factor, then we get

UAMP = [(10.1 m/s) (-0.5)]

UAMP = - 5.05 m/s